# Self-distributive structures in physics

### 21 Mar 2024 v1

@@ Abstract.
It is an important feature of our existing physical theories that observables generate one-parameter groups of transformations. In classical Hamiltonian mechanics and quantum mechanics, this is due to the fact that the observables form a Lie algebra, and it manifests itself in Noether’s theorem. In this paper, we introduce Lie quandles as the minimal mathematical structure needed to express the idea that observables generate transformations. This is based on the notion of a quandle used primarily in knot theory, whose main defining property is the self-distributivity equation $x\triangleright(y\triangleright z)=(x\triangleright y)\triangleright(x\triangleright z)$. We argue that Lie quandles can be thought of as nonlinear generalizations of Lie algebras.
@@
Acknowledgements. We thank Benjamin Walder for helpful discussions on Lie quandles.
@@ 1. Introduction
Observables play a dual role in our existing physical theories. On the one hand, they are the quantities that we think of as being measurable in principle. On the other hand, they are also exactly those quantities that generate groups of transformations, including symmetries and the dynamics of a theory. This surprising dual role is especially apparent and well-understood in Hamiltonian mechanics, where it is closely tied to the symplectic structure of phase space, and in quantum mechanics, where it manifests itself in the correspondence between self-adjoint operators and one-parameter groups of unitaries (Stone’s theorem). In both cases, this duality is what underlies Noether’s theorem: an observable is a conserved quantity if and only if the transformation group that it generates is a group of symmetries.
@@
This intriguingly self-referential equation is a form of self-distributivity, which famously appears in the theory of self-distributive structures like quandles [3, 4]. Surprisingly, the role of such self-distributivity in physics does not seem to have been appreciated before [^1] One notable exception is a blog post by Daniel Moskovich [5], which expresses the idea that associative algebraic structures are geometric in nature, while self-distributive operations are appropriate for ‘information physics’. However, he does not seem to have recognized the role of self-distributivity in the correspondence between observables and transformation groups., and quandles have hitherto mainly been considered in the context of knot theory. Hence the main purpose of this paper is to point out the relevance of self-distributivity in physics.
There is another appearance of self-distributivity in physics which underlines its relevance, but now at the level of states rather than observables. In quantum theory, forming convex combinations or mixtures of (mixed) states has been recognized as a crucial piece of structure already by von Neumann [6, Ch. IV], and has since become central through the framework of general probabilistic theories. Now if we consider convex combinations operations in the form
@@
    The thesis of Crans [7, Section 3.2.4] considers unital spindles and quandles internal to the category of coalgebras, and uses these structures as an intermediate notion between Lie groups and Lie algebras. Since the underlying set of her quandles correspond to the symmetric power of a Lie algebra, this seems quite different from our Lie quandles (of which the simplest examples are Lie algebras themselves).
@@
    Smooth quandles were introduced by Ishikawa222ishikawa; we also refer to the thesis of Yonemura [8] for a detailed exposition. Smooth quandles are quandles internal to the category of manifolds (meaning that the underlying set is a manifold and the quandle operation is smooth). This is more closely related to Lie quandles: a Lie quandle becomes a smooth quandle if one forgets all operations $\triangleright_{t}$ for $t\neq 1$.
@@ 2. Self-distributivity
Here we have adopted the terminology of Crans [7, Definition 52] [^3] We the attribute ‘left’ since we will not consider right shelves.. A good way to think about this equation is as saying that the map $x\triangleright-:S\to S$ is a homomorphism with respect to $\triangleright$ itself for every $x\in S$. In this way, shelves axiomatize the idea of a set acting on ‘itshelf’.
@@
may be the historically first example of a shelf [9, p. 481]. Note that every element acts by permuting the other two.
@@
    Similarly, we can work with points on the sphere $S^{2}$ and define $x\triangleright y$ to be the point obtained by rotating $y$ around $x$ by $\theta$ [10, Definition 4.3].
@@
    We can also take $S$ to be any symmetric space, such as the sphere $S^{2}$, and define $x\triangleright y$ to be the point obtained by reflecting $y$ across $x$ [4, Example 56] [^5] Note that [4] considers right self-distributivity rather than the left version of Definition 2.1, so that their $x\triangleright y$ corresponds to our $y\triangleright x$..
@@
The following example is a simple variation on the standard Alexander quandle from the theory of knot invariants [4, Example 67].
@@
Moreover, it can be shown that $\triangleright_{-1}$ equips the original set $Q$ with another quandle structure, and $\triangleright_{1}$ and $\triangleright_{-1}$ also mutually distribute over each other [4, p. 95]. These statements are observations that naturally lead to the Lie quandles of Section 3 if one considers an entire family of operations $\triangleright_{t}$ indexed by $t\in\mathbb{R}$.
@@
    This construction works the same way if $G$ is replaced by any quandle and the action on $X$ by a quandle action in the sense of [11, § 3.2].
Hitherto, the theory of quandles has been primarily applied to knot theory. We refer to the original paper of Joyce [3] and to the more recent textbook [4] for expositions of this application. [^6] A reader who knows about anyon physics may wonder whether anyons provide another way in which quandles appear in physics. As far as we can see this is not the case, since anyons are modelled by braided fusion categories, and these do not seem to bear any direct relation to quandles.
@@ A Lie quandle is a smooth manifold $Q$ together with an operation
for any matrix $Y$ and any invertible $A$ [12, Proposition 2.3(6)]. We apply this with $A=e^{isX}$ in the third step of the following sequence of equations:
@@
The characterization (3.4) of the Lie quandle structure implies that every (real) Lie algebra $\mathfrak{g}\subseteq M_{n}(\mathbb{C})$ is a Lie quandle with respect to the operations $\triangleright_{t}$ restricted from all matrices to $\mathfrak{g}$. By Ado’s theorem, we can embed every Lie algebra into $M_{n}(\mathbb{C})$ for some $n$, and therefore every Lie algebra is a Lie quandle in a canonical way. If $G$ is any Lie group with Lie algebra $\mathfrak{g}$, then this Lie quandle structure on $\mathfrak{g}$ can also be written in terms of the exponential map $\exp:\mathfrak{g}\to G$ and the adjoint action of $G$ on $\mathfrak{g}$ [12, §4.3.3] as
@@
The correspondence between observables and one-parameter groups of transformations plays a crucial role in Noether’s theorem. [^9] We refer to [13, Theorem 5.58] for a general version of Noether’s theorem in classical field theory. How can we understand this result in the framework of Lie quandles?

### 10 Apr 2024 v2

@@ Abstract.
It is an important feature of our existing physical theories that observables generate one-parameter groups of transformations. In classical Hamiltonian mechanics and quantum mechanics, this is due to the fact that the observables form a Lie algebra, and it manifests itself in Noether’s theorem. In this paper, we introduce Lie quandles as the minimal mathematical structure needed to express the idea that observables generate transformations. This is based on the notion of a quandle used most famously in knot theory, whose main defining property is the self-distributivity equation $x\triangleright(y\triangleright z)=(x\triangleright y)\triangleright(x\triangleright z)$. We argue that Lie quandles can be thought of as nonlinear generalizations of Lie algebras.
@@
Acknowledgements. We thank Benjamin Walder for helpful discussions on Lie quandles and Markus Szymik for valuable pointers to the literature.
@@ 1. Introduction
Observables play a dual role in our existing physical theories [1]. On the one hand, they are the quantities that we think of as being measurable in principle. On the other hand, they are also exactly those quantities that generate groups of transformations, including symmetries and the dynamics of a theory. This surprising dual role is especially apparent and well-understood in Hamiltonian mechanics, where it is closely tied to the symplectic structure of phase space, and in quantum mechanics, where it manifests itself in the correspondence between self-adjoint operators and one-parameter groups of unitaries (Stone’s theorem). In both cases, this duality is what underlies Noether’s theorem: an observable is a conserved quantity if and only if the transformation group that it generates is a group of symmetries.
@@
This intriguingly self-referential equation is a form of self-distributivity, which famously appears in the theory of self-distributive structures like quandles [4, 5]. Surprisingly, the role of such self-distributivity in physics does not seem to have been appreciated before [^1] One notable exception is a blog post by Daniel Moskovich [6], which expresses the idea that associative algebraic structures are geometric in nature, while self-distributive operations are appropriate for ‘information physics’. However, he does not seem to have recognized the role of self-distributivity in the correspondence between observables and transformation groups., and quandles have hitherto mainly been considered in pure mathematics [^2] Their most well-known application is to knot theory [7, 4, 8]. However, the earliest appearance of (involutive) quandles seems to be as algebraic structures abstracting reflections [9], and they have also been used in number theory [10] and real algebra [11]. Brieskorn’s early survey [12] discusses relations to braid groups, root systems and singularity theory.. Hence the main purpose of this paper is to point out the relevance of self-distributivity in physics.
There is another appearance of self-distributivity in physics which underlines its relevance, but now at the level of states rather than observables. In quantum theory, forming convex combinations or mixtures of (mixed) states has been recognized as a crucial piece of structure already by von Neumann [13, Ch. IV], and has since become central through the framework of general probabilistic theories. Now if we consider convex combinations operations in the form
@@
    The thesis of Crans [14, Section 3.2.4] considers unital spindles and quandles internal to the category of coalgebras, and uses these structures as an intermediate notion between Lie groups and Lie algebras. Since the underlying set of her quandles correspond to the symmetric power of a Lie algebra, this seems quite different from our Lie quandles (of which the simplest examples are Lie algebras themselves).
@@
    Smooth quandles were introduced by Ishikawa [15] [^3] There is a more general notion of quandles in categories with finite products [16], which yields smooth quandles upon taking the category to be the category of manifolds.; we also refer to the thesis of Yonemura [17] for a detailed exposition. Smooth quandles are quandles internal to the category of manifolds (meaning that the underlying set is a manifold and the quandle operation is smooth). This is more closely related to Lie quandles: a Lie quandle becomes a smooth quandle if one forgets all operations $\triangleright_{t}$ for $t\neq 1$.
@@ 2. Self-distributivity
Here we have adopted the terminology of Crans [14, Definition 52] [^4] We the attribute ‘left’ since we will not consider right shelves.. A good way to think about this equation is as saying that the map $x\triangleright-:S\to S$ is a homomorphism with respect to $\triangleright$ itself for every $x\in S$. In this way, shelves axiomatize the idea of a set acting on ‘itshelf’.
@@
may be the historically first example of a shelf [18, p. 481]. Note that every element acts by permuting the other two.
@@
    Similarly, we can work with points on the sphere $S^{2}$ and define $x\triangleright y$ to be the point obtained by rotating $y$ around $x$ by $\theta$ [19, Definition 4.3].
@@
    We can also take $S$ to be any symmetric space, such as the sphere $S^{2}$, and define $x\triangleright y$ to be the point obtained by reflecting $y$ across $x$ [5, Example 56] [^6] Note that [5] considers right self-distributivity rather than the left version of Definition 2.1, so that their $x\triangleright y$ corresponds to our $y\triangleright x$..
@@
The following example is a simple variation on the standard Alexander quandle from the theory of knot invariants [5, Example 67]. It also appears in Moskovich’s blog post [6] and has probably been known before.
@@
Moreover, it can be shown that $\triangleright_{-1}$ equips the original set $Q$ with another quandle structure, and $\triangleright_{1}$ and $\triangleright_{-1}$ also mutually distribute over each other [5, p. 95]. These statements are observations that naturally lead to the Lie quandles of Section 3 if one considers an entire family of operations $\triangleright_{t}$ indexed by $t\in\mathbb{R}$.
@@
    This construction works the same way if $G$ is replaced by any quandle and the action on $X$ by a quandle action in the sense of [20, § 3.2].
Hitherto, the theory of quandles has been primarily applied to knot theory. We refer to the original papers [7, 4, 8] as well as to the more recent textbook [5] for expositions of this application. [^7] A reader who knows about anyon physics may wonder whether anyons provide another way in which quandles appear in physics. As far as we can see this is not the case, since anyons are modelled by braided fusion categories, and these do not seem to bear any direct relation (with physical significance) to quandles.
@@ A Lie quandle is a smooth manifold $Q$ together with an operation
for any matrix $Y$ and any invertible $A$ [21, Proposition 2.3(6)]. We apply this with $A=e^{isX}$ in the third step of the following sequence of equations:
@@
The characterization (3.4) of the Lie quandle structure implies that every (real) Lie algebra $\mathfrak{g}\subseteq M_{n}(\mathbb{C})$ is a Lie quandle with respect to the operations $\triangleright_{t}$ restricted from all matrices to $\mathfrak{g}$. By Ado’s theorem, we can embed every Lie algebra into $M_{n}(\mathbb{C})$ for some $n$, and therefore every Lie algebra is a Lie quandle in a canonical way. If $G$ is any Lie group with Lie algebra $\mathfrak{g}$, then this Lie quandle structure on $\mathfrak{g}$ can also be written in terms of the exponential map $\exp:\mathfrak{g}\to G$ and the adjoint action of $G$ on $\mathfrak{g}$ [21, §4.3.3] as
@@
The correspondence between observables and one-parameter groups of transformations plays a crucial role in Noether’s theorem. [^10] We refer to [22, Theorem 5.58] for a general version of Noether’s theorem in classical field theory. We can now understand this result by phrasing a formulation of Noether’s theorem due to Baez [1] in the framework of Lie quandles.